# Axial Neural Networks for Dimension-Free Foundation Models

### v1

@@ 1 Introduction
This paper is concerned with developing key techniques for building successful foundation models for physics data, such as climate time-series data and solutions of partial differential equations (PDEs). When training such a model, we often have to combine datasets from multiple systems or differential equations that operate on different dimensionalities. For instance, the Burgers equation describing dissipative fluid flow is usually studied in one spatial dimension, whereas the Navier–Stokes equations are studied in two or three dimensions. Since the solutions of these equations are typically represented as tensors whose elements are points in spatial and temporal grids, the different dimensionalities mean that those tensors storing solutions have different numbers of axes.
@@ 2.2 Transpose and Axis-Permutation Equivariance
$\text{T}_{ij}(\boldsymbol{x})_{d_{1}\cdots d_{i}\cdots d_{j}\cdots d_{K}}=x_{d_{1}\cdots d_{j}\cdots d_{i}\cdots d_{K}},$ (3)
where $\text{T}_{ij}$ denotes the transpose between axes $i$ and $j$. An axis permutation is obtained by cumulative transposes, corresponding to reordering the axes. For a rank-4 tensor $\boldsymbol{x}$, for example,
@@ 3.1 Set-Based Axial Neural Networks
Let $\phi$ be an axis-permutation equivariant function (e.g., another SXNN or pointwise operation). The SXNN in Eq.˜8 is axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{x}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$.
Proof.
§˜B.1. ∎
@@
where $\operatorname{Linear}_{3}$ applies a linear layer along the third axis ($D$), $\operatorname{Pool}_{1,2}$ pools over the other two axes ($H$ and $W$), and they treat the remaining axis as batch dimensions. Note that only three out of six permutations are required, since pooling is invariant to permutations across the pooled dimensions, i.e. $\text{Pool}(\mathbb{R}^{H\!\times\!W\!\times\!D})=\text{Pool}(\mathbb{R}^{W\!\times\!H\!\times\!D})$, which omits the redundant permutations. Note also that pooling is equivalent to the approach commonly used in 3D inflation of 2D convolutional layers [6, 28]. The transformation flow of the feature sizes in Eq.˜9 is summarized as
@@
Non-linearity. $\phi$ and $\psi$ in Eq.˜8 can be arbitrary neural networks, including those with a single non-linearity such as ReLU or Sigmoid. Therefore, using any type of pointwise operation (including non-linearities) does not violate the axis-permutation equivariance of XNN.
@@
where * denotes convolution, the first matrix is the kernel, and the second is a $2\!\times\!2$ patch from the input image. On the other hand, the axial convolution in Eq.˜11 with average pooling results in:
@@ 3.2 Graph-Based Axial Neural Networks
We imagine an undirected graph over the axes such as Fig.˜1. We now apply the message-passing logic of GNNs as described in Eq.˜1. Since the neighbors of $H$ are $W$ and $D$ and likewise for the other axes, we can informally express the update rule as
$\boldsymbol{h}_{H}$ $=\phi\left(H,\psi(H,H)\oplus\psi(H,W)\oplus\psi(H,D)\right),$ $\boldsymbol{h}_{W}$ $=\phi\left(W,\psi(W,W)\oplus\psi(W,D)\oplus\psi(W,H)\right),$
$\boldsymbol{h}_{D}$ $=\phi\left(D,\psi(D,D)\oplus\psi(D,H)\oplus\psi(D,W)\right),$ (14)
where $\oplus$ denotes an arithmetic form of the permutation-invariant operator $\bigoplus$ and $H,W,D$ are used as a conceptual feature representing each axis. The functions $\phi$ and $\psi$ can be chosen based on the architecture or task.
For example, in the case of convolutional layers, let $\boldsymbol{x}\in\mathbb{R}^{H\!\times\!W\!\times\!D\!\times\!C}$, where $C$ is the number of channels. We can define $\phi$ as the identity function that returns the second argument (i.e. $\phi(A,B)=B$) and $\psi(H,W)$ as a $\operatorname{Conv2D}$ applied over axes $H$ and $W$ with the pooling layer at the end for matching the tensor sizes. We omit $\psi(H,H)$ as it is nontrivial, and the absence of it does not violate the axis permutation equivariance. Then §˜3.2 becomes
$\boldsymbol{h}_{H}$ $=\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\Pi_{4}(\boldsymbol{x})+\text{T}_{23}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\Pi_{5}(\boldsymbol{x}),$ (15)
$\boldsymbol{h}_{W}$ $=\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\Pi_{2}(\boldsymbol{x})+\text{T}_{23}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\Pi_{3}(\boldsymbol{x}),$
$\boldsymbol{h}_{D}$ $=\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\Pi_{0}(\boldsymbol{x})+\text{T}_{23}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\Pi_{1}(\boldsymbol{x}).$
Of course, $\phi$ and $\psi$ need not be linear, and it can be a multilayer perceptron. The transformation flow of the feature sizes of $\boldsymbol{h}_{H}$ in Eq.˜15 would be:
$H\!\times\!W\!\times\!D\begin{Bmatrix}\xrightarrow{\Pi_{4}}D\!\times\!\boldsymbol{W}\!\times\!\boldsymbol{H}\xrightarrow{\text{Conv2D}_{2,3}}\boldsymbol{D}\!\times\!w\!\times\!h\xrightarrow{\text{Pool}_{1}}d\!\times\!w\!\times\!h\xrightarrow{}\\ \xrightarrow{\Pi_{5}}W\!\times\!\boldsymbol{D}\!\times\!\boldsymbol{H}\xrightarrow{\text{Conv2D}_{2,3}}\boldsymbol{W}\!\times\!d\!\times\!h\xrightarrow{\text{Pool}_{1}}w\!\times\!d\!\times\!h\xrightarrow{\text{T}_{23}}\end{Bmatrix}\xrightarrow{+}d\!\times\!w\!\times\!h=\boldsymbol{h}_{H}.$
Here the channel axis $C$ is omitted for simplicity. The convolution example is also illustrated in Fig.˜2(a).
For 1D and 2D cases, §˜3.2 reduces to generating the corresponding number of outputs.
@@
The 1D case does not consider interaction with the other nodes and $\psi(H,H)$ is still nontrivial. Thus, instead of determining $\psi(H,H)$, we omit the self-edge term $\psi(H,H)$ or $\psi(W,W)$ but rather augment a 1D tensor to a 2D tensor to fully utilize $\operatorname{Conv2D}$. Possible augmentations include the outer product, repetition, and the diagonal matrix. We adopt repetition, which repeats the 1D tensor along a new axis to match the kernel size of Conv2D, and then averages it after lifting to recover the original 1D tensor. Generalization of the lifting layers, §˜3.2 and Eq.˜16, to rank-$K$ tensor is described as
$\operatorname{(\textbf{Lifting})}\quad\boldsymbol{h}_{d_{i}}=\phi\bigg(d_{i},\bigoplus_{j\neq i}^{K}\psi(d_{i},d_{j})\bigg),\quad K>1.$ (17)
@@
The lifting layer of GXNN, Eq.˜17, is axis-permutation equivariant for any rank-$K$ (except $K=1$) tensor $\boldsymbol{x}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$.
Proof.
§˜B.2. ∎
@@
$\boldsymbol{h}^{\prime}_{H}$ $=\phi\left(\boldsymbol{h}_{H},\psi(\boldsymbol{h}_{H},\boldsymbol{h}_{H})\oplus\psi(\boldsymbol{h}_{H},\boldsymbol{h}_{W})\oplus\psi(\boldsymbol{h}_{H},\boldsymbol{h}_{D})\right),$ (18)
$\boldsymbol{h}^{\prime}_{W}$ $=\phi\left(\boldsymbol{h}_{W},\psi(\boldsymbol{h}_{W},\boldsymbol{h}_{W})\oplus\psi(\boldsymbol{h}_{W},\boldsymbol{h}_{D})\oplus\psi(\boldsymbol{h}_{W},\boldsymbol{h}_{H})\right),$
$\boldsymbol{h}^{\prime}_{D}$ $=\phi\left(\boldsymbol{h}_{D},\psi(\boldsymbol{h}_{D},\boldsymbol{h}_{D})\oplus\psi(\boldsymbol{h}_{D},\boldsymbol{h}_{H})\oplus\psi(\boldsymbol{h}_{D},\boldsymbol{h}_{W})\right),$
Therefore, the subsequent layers should also be dimension-agnostic message-passing architectures, with the lifted features $\boldsymbol{h}_{H}$, $\boldsymbol{h}_{W}$, and $\boldsymbol{h}_{D}$ as graph nodes.
For example, in the case of self-attention layers, we can define $\phi$ as a residual path, i.e. $\phi(A,B(A))=A+B(A)$, and $\psi$ as a self-attention applied over the last axis of the second tensor, i.e. $\psi(A,B)=\operatorname{SelfAttn}_{K}(B)$. Then, as also described in Fig.˜2(b), the GNN in Eq.˜18 becomes
$\boldsymbol{h}^{\prime}_{H}$ $=\boldsymbol{h}_{H}+\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{H})+\Pi_{4}\Pi_{2}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{W})+\Pi_{4}\Pi_{0}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{D}),$ (19)
$\boldsymbol{h}^{\prime}_{W}$ $=\boldsymbol{h}_{W}+\Pi_{2}\Pi_{4}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{H})+\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{W})+\Pi_{2}\Pi_{0}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{D}),$
$\boldsymbol{h}^{\prime}_{D}$ $=\boldsymbol{h}_{D}+\Pi_{0}\Pi_{4}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{H})+\Pi_{0}\Pi_{2}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{W})+\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{D}),$
@@
$\boldsymbol{h}^{\prime}_{H}=\boldsymbol{h}_{H}+\Pi_{4}(\boldsymbol{m}),\quad\boldsymbol{h}^{\prime}_{W}=\boldsymbol{h}_{W}+\Pi_{2}(\boldsymbol{m}),\quad\boldsymbol{h}^{\prime}_{D}=\boldsymbol{h}_{D}+\Pi_{0}(\boldsymbol{m}),$
[…]
$\operatorname{(\textbf{Subsequent})}\quad\boldsymbol{h}^{\prime}_{d_{i}}=\phi\bigg(\boldsymbol{h}_{d_{i}},\bigoplus_{j=1}^{K}\psi(\boldsymbol{h}_{d_{i}},\boldsymbol{h}_{d_{j}})\bigg).$ (21)
@@
The subsequent layers of GXNN, Eq.˜21, are axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{h}_{d_{i}}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$.
Proof.
§˜B.3. ∎
@@
$(\textbf{rank }3)\quad\boldsymbol{h}^{\prime}$ $=\Pi_{4}^{-1}(\boldsymbol{h}_{H}^{\prime})\oplus\Pi_{2}^{-1}(\boldsymbol{h}_{W}^{\prime})\oplus\Pi_{0}^{-1}(\boldsymbol{h}_{D}^{\prime}),$ $(\textbf{rank }K)\quad\boldsymbol{h}^{\prime}$ $=\bigoplus_{i=1}^{K}\boldsymbol{h}^{\prime}_{d_{i}}.$ (22)

### v2

@@ 1 Introduction
This paper is concerned with developing key techniques for building successful foundation models for physics data, such as climate time-series data and solutions of partial differential equations (PDEs). When training such a model, we often have to combine datasets from multiple systems or differential equations that operate on different dimensionalities. For instance, the Burgers equation describing dissipative fluid flow is usually studied in one spatial dimension, whereas the Navier–Stokes equations are studied in two or three dimensions. Since the solutions of these equations are typically represented as tensors whose elements are points in spatial and temporal grids, different dimensionalities mean that those tensors storing solutions have different numbers of axes.
@@ 2.2 Transpose and Axis-Permutation Equivariance
$\boldsymbol{x}^{\top_{ij}}_{d_{1}\cdots d_{i}\cdots d_{j}\cdots d_{K}}=x_{d_{1}\cdots d_{j}\cdots d_{i}\cdots d_{K}},$ (3)
where $\top_{ij}$ denotes the transpose between axes $i$ and $j$. An axis permutation is obtained by cumulative transposes, corresponding to reordering the axes. For a rank-4 tensor $\boldsymbol{x}$, for example,
@@
2.3 Cycle Notation for Axes Permutation
To express the reordering of axes throughout this paper, we require a precise notation for permutations. A permutation is formally defined as a bijection (a one-to-one mapping) from a set onto itself. In our context, the set consists of axis indices, such as $\{H,W,D\}$ for $\mathbf{x}\in\mathbb{R}^{H\!\times\!W\!\times\!D}$.
First, we review the standard cycle notation common in algebra. This notation uses parentheses $()$ to group elements into disjoint cycles that show the path each element follows under the permutation. An element within a cycle is mapped to the element immediately following it. The last element in a cycle is mapped back to the first, completing the loop.
For example, consider permutations on the set of three axis indices $\{1,2,3\}$:
  • A permutation that maps $1\to 3$, $3\to 2$, and $2\to 1$ is written as the single cycle (1 3 2)(1\ 3\ 2).
  • A permutation that maps $1\to 3$, $3\to 1$, and leaves 2 unchanged, $2\to 2$, is written as (1 3)(1\ 3). The element 2 is a fixed point.
Adopting this cycle notation, we define the transformation $\text{T}_{i_{1}i_{2}\dots i_{n}}$ as the permutation that maps the original ordered set of axes $d_{1}\!\times\!d_{2}\!\times\!\cdots\!\times\!d_{n}$ to the new ordered arrangement $d_{i_{1}}\!\times\!d_{i_{2}}\!\times\!\cdots\!\times\!d_{i_{n}}$. For example,
$\boldsymbol{y}=\text{T}_{132}(\boldsymbol{x}),\quad\boldsymbol{y}\in\mathbb{R}^{W\!\times\!D\!\times\!H},\quad\boldsymbol{x}\in\mathbb{R}^{H\!\times\!W\!\times\!D},$ (7)
$\boldsymbol{y}^{\prime}=\text{T}_{13}(\boldsymbol{x}),\quad\boldsymbol{y}^{\prime}\in\mathbb{R}^{D\!\times\!W\!\times\!H},\quad\boldsymbol{x}\in\mathbb{R}^{H\!\times\!W\!\times\!D}.$
@@ 3.1 Set-Based Axial Neural Networks
Let $\phi$ be an axis-permutation equivariant function (e.g., another SXNN or pointwise operation). The SXNN in Eq.˜9 is axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{x}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$. (§˜B.1 for proof)
@@
where $\operatorname{Linear}_{3}$ applies a linear layer along the third axis ($D$), $\operatorname{Pool}_{1,2}$ pools over the other two axes ($H$ and $W$), and they treat the remaining axis as batch dimensions. Note that only three out of six permutations are required, since pooling is invariant to permutations across the pooled dimensions, i.e. $\text{Pool}(\mathbb{R}^{H\!\times\!W\!\times\!D})=\text{Pool}(\mathbb{R}^{W\!\times\!H\!\times\!D})$, which omits the redundant permutations. Note also that pooling is equivalent to the approach commonly used in 3D inflation of 2D convolutional layers [6, 28]. The transformation flow of the feature sizes in Eq.˜10 is summarized as
@@
Non-linearity. $\phi$ and $\psi$ in Eq.˜9 can be arbitrary neural networks, including those with a single non-linearity such as ReLU or Sigmoid. Therefore, using any type of pointwise operation (including non-linearities) does not violate the axis-permutation equivariance of XNN.
@@
where * denotes convolution, the first matrix is the kernel, and the second is a $2\!\times\!2$ patch from the input image. On the other hand, the axial convolution in Eq.˜12 with average pooling results in:
@@ 3.2 Graph-Based Axial Neural Networks
We imagine an undirected graph over the axes such as Fig.˜1. We now apply the message-passing logic of GNNs as described in Eq.˜1. Since the neighbors of $d_{1}$ are $d_{2},\ldots,d_{K}$ and likewise for the other axes, we can informally express the update rule as
$\boldsymbol{h}_{d_{i}}=\phi\bigg(d_{i},\bigoplus_{j=1}^{K}\psi(d_{i},d_{j})\bigg)$ (15)
where $\oplus$ denotes an arithmetic form of the permutation-invariant operator $\bigoplus$ and $d_{1},d_{2},\ldots,d_{K}$ are used as a conceptual feature representing each axis. The functions $\phi$ and $\psi$ can be chosen based on the architecture or task. Although Eq.˜15 follows parameter-sharing principles and permutation equivariance, in practice, we have to modify it to match the axes order after the aggregation $\oplus$.
For example, in the case of convolutional layers, let $\boldsymbol{x}\in\mathbb{R}^{H\!\times\!W\!\times\!D\!\times\!C}$, where $C$ is the number of channels. We can define $\phi$ as the identity function that returns the second argument (i.e. $\phi(A,B)=B$) and $\psi(H,W)$ as a $\operatorname{Conv2D}$ applied over axes $H$ and $W$ with the pooling layer at the end for matching the tensor sizes. We omit $\psi(H,H)$ as it is nontrivial, and the absence of it does not violate the axis permutation equivariance. When the indices of $\{H,W,D\}$ are $\{1,2,3\}$, Eq.˜15 becomes
$\boldsymbol{h}_{H}$ $=\text{T}_{13}\text{T}_{13}^{-1}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\text{T}_{13}(\boldsymbol{x})+\text{T}_{13}\text{T}_{132}^{-1}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\text{T}_{132}(\boldsymbol{x}),$ (16)
$\boldsymbol{h}_{W}$ $=\text{T}_{23}\text{T}_{23}^{-1}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\text{T}_{23}(\boldsymbol{x})+\text{T}_{23}\text{T}_{123}^{-1}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\text{T}_{123}(\boldsymbol{x}),$
$\boldsymbol{h}_{D}$ $=\text{T}_{33}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}(\boldsymbol{x})+\text{T}_{33}\text{T}_{12}^{-1}\operatorname{Pool}_{1}\operatorname{Conv2D}_{2,3}\text{T}_{12}(\boldsymbol{x}),$
where $\text{T}_{ijk}$ denotes reordering axes $1,2,3$ to $i,j,k$ as explained in §˜2.3. Of course, $\phi$ and $\psi$ need not be linear, and it can be a multilayer perceptron. The transformation flow of the feature sizes of $\boldsymbol{h}_{H}$ in Eq.˜16 would be:
$H\!\times\!W\!\times\!D\begin{Bmatrix}\xrightarrow{\text{T}_{13}}D\!\times\!\boldsymbol{W}\!\times\!\boldsymbol{H}\xrightarrow{\text{Conv2D}_{2,3}}\boldsymbol{D}\!\times\!w\!\times\!h\xrightarrow{\text{Pool}_{1}}d\!\times\!w\!\times\!h\xrightarrow{\text{T}_{13}\text{T}_{13}^{-1}}\\ \xrightarrow{\text{T}_{132}}W\!\times\!\boldsymbol{D}\!\times\!\boldsymbol{H}\xrightarrow{\text{Conv2D}_{2,3}}\boldsymbol{W}\!\times\!d\!\times\!h\xrightarrow{\text{Pool}_{1}}w\!\times\!d\!\times\!h\xrightarrow{\text{T}_{13}\text{T}_{132}^{-1}}\end{Bmatrix}\xrightarrow{+}d\!\times\!w\!\times\!h=\boldsymbol{h}_{H}.$
Here the channel axis $C$ is omitted for simplicity. A CNN example is also illustrated in Fig.˜2(a).
For 1D and 2D cases, Eq.˜15 reduces to generating the corresponding number of outputs.
@@
The 1D case does not consider interaction with the other nodes and $\psi(H,H)$ is still nontrivial. Thus, instead of determining $\psi(H,H)$, we omit the self-edge term $\psi(H,H)$ or $\psi(W,W)$ but rather augment a 1D tensor to a 2D tensor to fully utilize $\operatorname{Conv2D}$. Possible augmentations include the outer product, repetition, and the diagonal matrix. We adopt repetition, which repeats the 1D tensor along a new axis to match the kernel size of Conv2D, and then averages it after lifting to recover the original 1D tensor.
Generalization of the $\operatorname{\textbf{Lifting}}$ layers, Eq.˜16, to rank-$K$ tensor is described as
$\boldsymbol{h}_{d_{i}}=\phi\bigg(\text{T}_{(i)(K)}(\boldsymbol{x}),\bigoplus_{j\neq i}^{K}\text{T}_{(i)(K)}\text{T}_{(j)(K-1)(i)(K)}^{-1}\psi\left(\text{T}_{(j)(K-1)(i)(K)}(\boldsymbol{x})\right)\bigg),\quad K>1,$ (18)
where $\text{T}_{(a)(b)(c)(d)}=\text{T}_{abcd}$.
@@
Under some assumptions, the lifting layer of GXNN, Eq.˜18, is axis-permutation equivariant for any rank-$K$ (except $K=1$) tensor $\boldsymbol{x}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$. (§˜B.2 for assumptions and proof)
@@
$\boldsymbol{h}^{\prime}_{d_{i}}=\phi\bigg(\boldsymbol{h}_{d_{i}},\bigoplus_{j=1}^{K}\psi(\boldsymbol{h}_{d_{i}},\boldsymbol{h}_{d_{j}})\bigg).$ (19)
Therefore, the subsequent layers should also be dimension-agnostic message-passing architectures, with the lifted features $\boldsymbol{h}_{H}$, $\boldsymbol{h}_{W}$, and $\boldsymbol{h}_{D}$ as graph nodes. Eq.˜19 also needs to be modified to match the indices order in the aggregation.
For example, in the case of self-attention layers, we can define $\phi$ as a residual path, i.e. $\phi(A,B(A))=A+B(A)$, and $\psi$ as a self-attention applied over the last axis of the second tensor, i.e. $\psi(A,B)=\operatorname{SelfAttn}_{K}(B)$. Then, as also described in Fig.˜2(b), the GNN in Eq.˜19 becomes
$\boldsymbol{h}^{\prime}_{H}=\boldsymbol{h}_{H}+\text{T}_{13}\text{T}_{13}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{H})+\text{T}_{13}\text{T}_{23}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{W})+\text{T}_{13}\text{T}_{33}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{D}),$
$\boldsymbol{h}^{\prime}_{W}=\boldsymbol{h}_{W}+\text{T}_{23}\text{T}_{13}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{H})+\text{T}_{23}\text{T}_{23}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{W})+\text{T}_{23}\text{T}_{33}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{D}),$
$\boldsymbol{h}^{\prime}_{D}=\boldsymbol{h}_{D}+\text{T}_{33}\text{T}_{13}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{H})+\text{T}_{33}\text{T}_{23}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{W})+\text{T}_{33}\text{T}_{33}^{-1}\operatorname{SelfAttn}_{3}(\boldsymbol{h}_{D}),$ (20)
@@
$\boldsymbol{h}^{\prime}_{H}=\boldsymbol{h}_{H}+\text{T}_{13}(\boldsymbol{m}),\quad\boldsymbol{h}^{\prime}_{W}=\boldsymbol{h}_{W}+\text{T}_{23}(\boldsymbol{m}),\quad\boldsymbol{h}^{\prime}_{D}=\boldsymbol{h}_{D}+\text{T}_{33}(\boldsymbol{m}),$
where Ts are used for aligning the axis order to match the axes in the aggregation. Similarly, the $\operatorname{\textbf{Subsequent}}$ layers in §˜3.2 for a rank-$K$ tensor can be written as
$\boldsymbol{h}^{\prime}_{d_{i}}=\phi\bigg(\boldsymbol{h}_{d_{i}},\text{T}_{(i)(K)}\bigoplus_{j=1}^{K}\text{T}_{(j)(K)}^{-1}\psi(\boldsymbol{h}_{d_{j}})\bigg).$ (22)
@@
Under some assumptions, the subsequent layers of GXNN in Eq.˜22 are axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{h}_{d_{i}}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$. (§˜B.3 for assumptions and proof)
@@
$(\textbf{rank }3)\quad\boldsymbol{h}^{\prime}$ $=\text{T}_{13}^{-1}(\boldsymbol{h}_{H}^{\prime})\oplus\text{T}_{23}^{-1}(\boldsymbol{h}_{W}^{\prime})\oplus\text{T}_{33}^{-1}(\boldsymbol{h}_{D}^{\prime}),$ $(\textbf{rank }K)\quad\boldsymbol{h}^{\prime}$ $=\bigoplus_{i=1}^{K}\text{T}_{(i)(K)}^{-1}\boldsymbol{h}^{\prime}_{d_{i}}.$ (23)